Pls. help me
The longer side of a rectangle is 4 less than twice the shorter side. The perimeter of the rectangle is 40. Find the length of each side of the rectangle.

Answer by stanbon(75887)

The longer side of a rectangle is 4 less than twice the shorter side. The perimeter of the rectangle is 40. Find the length of each side of the rectangle.
-------------------------------
Let the shorter side be "x".
The longer side is "2x-4".
-------------------------------
Perimeter = 2(shorter + longer)
40 = 2(x + 2x-4)
20 = 3x-4
3x = 24
x = 8 units (shorter side; there are 2 "shorter" sides)
2x-4 = 16-4 = 12 (longer side; there are 2 "longer sides)
---
Cheers,
Stan H.
---